OpenStudy (pphalke):

The US Senate has 100 members. After an certain election there were 8 more Democrats than Republicans with no other parties represented. How many members of each party were there in the senate?

OpenStudy (dayakar):

let no of republicans = x no. of Democrats = x+8 [ 8 more than republicans] total = 100 x+x+8 =100 can u find x now?

OpenStudy (pphalke):

10

OpenStudy (dayakar):

do u understand

OpenStudy (dayakar):

how do u get it?

OpenStudy (pphalke):

10 is the number of democrats? what about republicans

OpenStudy (dayakar):

x+x+8 =100 2x+8 =100

OpenStudy (dayakar):

how?

OpenStudy (dayakar):

2x+8 =100 2x= 100-8 2x= 92 x= 92/2 x=46 no .of republican members = x = 46 rest of them are democrats